hello , can you please help me on this equation .
Find the perimeter and area of the rectangle with length of 8 sqrt (7) inches and width of 2 sqrt (7) - 3 sqrt(5) inches.

Answer by stanbon(75887)

Find the perimeter and area of the rectangle with
length of 8 sqrt (7) inches and
width of 2 sqrt (7) - 3 sqrt(5) inches.
-------------------------
Perimeter = 2(length + width)
P = 2(8sqrt(7) + 2sqrt(7)-3)
---
P = 2(10sqrt(7) - 3)
P = [20sqrt(7) -6] inches
=================================
Area = length * width
A = (8sqrt(7)*(2sqrt(7)-3))
---
A = 16*7 - 24sqrt(7)
---
Area = 112 - 24sqrt(7)
==========================
Cheers,
Stan H.
===========